OpenStudy (anonymous):

11y^3/10p^6 x 100p^8/121y

OpenStudy (anonymous):

hi ms juanita how r u?

OpenStudy (anonymous):

right ???? 10y^2p^2 / 11

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

i really enjoy helping others with my fav subject. keep them coming

OpenStudy (anonymous):

is the 2/11 in fraction form or to the power

OpenStudy (accessdenied):

\[ \begin{split} \frac{11y^{3}}{10p^{6}} \frac{100p^{8}}{121y} &= \frac{11*100 p^{8}y^{3}}{121*10 p^{6}y}\\ &= \frac{1100}{1210}p^{8-6}y^{3-1}\\ &= \frac{110}{121}p^{2}y^{2}\\ &= \frac{10}{11}y^{2}p^{2}\\ \end{split} \]